OpenStudy (acespeedfighter):

Bill and Lisa are surveying their classmates about their summer reading. Their questions are given below: Bill: How many books were recommended by the book reading club this month? Lisa: How many books did you read from the recommended list? Who wrote a statistical question and why? (5 points) Lisa, because there can be only one answer to the question Bill, because there can be many answers to the question Lisa, because there will be variability in the responses collected Bill, because every student will give the same answer

OpenStudy (acespeedfighter):

KYLE OMG I THOUGH I LOST DUDE

OpenStudy (kgrendel0324):

Haha nah im still alive

OpenStudy (acespeedfighter):

(mumbles) hmm were could he have gone

OpenStudy (acespeedfighter):

so...kyle can you help :D

OpenStudy (kgrendel0324):

Yeah sure

OpenStudy (acespeedfighter):

Viper?

OpenStudy (kgrendel0324):

Yup

OpenStudy (acespeedfighter):

hmm

OpenStudy (acespeedfighter):

I think its D but im not to sure if it is

OpenStudy (kgrendel0324):

Okay, you have to ask yourself while looking at both of them. Will this one have the same answer?

OpenStudy (kgrendel0324):

If that is yes, then it is a statistic, if not, then it isn't

OpenStudy (kgrendel0324):

Which answer choices can you eliminate knowing the above information?

OpenStudy (acespeedfighter):

so,,,,,again with Albert Einstein theory Answer = Medal

OpenStudy (acespeedfighter):

or A = M^2

OpenStudy (acespeedfighter):

You heard of it right?

OpenStudy (kgrendel0324):

mc^2

OpenStudy (acespeedfighter):

Hmmm i dont think so...

OpenStudy (kgrendel0324):

It is D to reinforce your first theory

OpenStudy (acespeedfighter):

Yep Eisteins theory helped diddnt it

OpenStudy (kgrendel0324):

Yeah Albert Einstein once said, E=mc^2

OpenStudy (acespeedfighter):

A = M^2*